OpenStudy (anonymous):

Is this equation right for this? On a class trip, there were 45 more girls than boys. The total number of students on the trip was 211

OpenStudy (anonymous):

221 = (x+45)x?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

x + (x+45)

OpenStudy (anonymous):

if x = number ofboys

OpenStudy (lilg132):

TS = X + 45

OpenStudy (lilg132):

ts = total students x = boys

OpenStudy (anonymous):

lilg - that would be wrong. TS = x + x + 45 (boys plus girls and girls are 45 more than boys). => 211 = 2x + 45